calculate the mean median from the following data and find out mode using relation among mean,median and mode. Class; 0-10 10-20 20-30 30-40 40-50 50-60 60-70 Frequency; 10 20 30 40 50 60 70


1
Expert's answer
2022-01-03T09:23:13-0500

1.Mean

We create a table by adding the columns to given data. The first column is for midpoint of class, "x_i"

and other column for product of frequency and midpoint of class "f_ix_i"

as shown below.




"\bar x=\frac{∑f_ix_i}{∑f_i}"

"\bar x=\frac{25+150+450+1050+900+660+325}{5+10+18+30+20+12+5}\\\bar x=35.6"

b)Median

We create a table by adding the columns to given data. First column is for cumulative frequency which is sum of frequencies up to that class interval as follows;




"median=L+(\frac{{\frac{N}{2}}−f_0}{f})h"

"median=30+(\frac{{\frac{100}{2}}−33}{30})10"

"=35.67"


c.(mode)




"Mode=L+(\frac{f_1−f_0}{2f_1−f_0−f_2})h"


"Mode=30+(\frac{30−18}{2×30−18−20})×10"

"=35.45"